How much charge passes in 2 minutes through a junction through which a steady current of 5A exists?
1
Expert's answer
2012-04-03T11:59:02-0400
The coulomb is defined as the charge transported by a steady current of one ampere in one second. So,

Q = I·T = 5*120 = 600 C.